OpenStudy (anonymous):

\[f(x)=ln(1+x)\] \[ln(1+x)=\frac{d}{dx}(\frac{1}{1-x})\] \[=\frac{d}{dx} \sum_{n=0}^{\infty} x^n\] \[=\sum_{n=0}^{\infty}nx^{n-1}\] \[|x|<1\] -1

OpenStudy (anonymous):

Please check, I think this is correct

OpenStudy (anonymous):

Sorry , there is smthing wrong with your answer !

OpenStudy (anonymous):

The problem reads "Please find a power series representation", and find the radius of convergence.

OpenStudy (anonymous):

R=1

OpenStudy (anonymous):

yeah ! that's right ! but in the first line...look at what you wrote !

OpenStudy (anonymous):

so I guess I did it wrong because the function is \[f(x)= ln(1+x)\]

OpenStudy (anonymous):

sorry !!the second !!

OpenStudy (anonymous):

so that would make it \[=\frac{d}{dx} \sum_{n=0}^{\infty} (-1)^nx^n\]

OpenStudy (anonymous):

and \[=\sum_{n=0}^{\infty}(-1)^nnx^{n-1}\]

OpenStudy (anonymous):

from 1 to infinity....But what you wrote still correct !

OpenStudy (anonymous):

why is it from 1 to infinity instead of zero?

OpenStudy (anonymous):

*instead of zero to infinity?

OpenStudy (anonymous):

\[ \log (x+1)=\sum _{n=1}^{\infty } \frac{(-1)^{n+1} x^n}{n} \]

OpenStudy (anonymous):

ok now I'm really confused

OpenStudy (anonymous):

\[ \log (x+1)=\int_0^x \frac{1}{t+1} \, dt \]

OpenStudy (anonymous):

I'm going to sleep ! I didn't pay attention to that ! this is what I was telling you from the start ! \[d(\ln(1+x))/dx)=1/(1+x)\] not what you wrote !

OpenStudy (anonymous):

You have to integrate the series of \(\frac 1 {1+t}\) not differentiating it.

OpenStudy (anonymous):

oh I see...

OpenStudy (anonymous):

Did you get it?

OpenStudy (anonymous):

what was your thought process from \[\int_0^x\frac 1 {1+t}dt\] to \[\sum_{n=0}^\infty (-1)^{n+1} \frac {x^{n+1} } { n+1 }\\\]

OpenStudy (anonymous):

Ohhh....I forgot to read the last section on Paul's online notes http://tutorial.math.lamar.edu/Classes/CalcII/PowerSeriesandFunctions.aspx

OpenStudy (anonymous):

@eliassaab \[1/(1+t)=\sum_{0}^{\infty} (-1)^nt^n\] isn't like this ?

OpenStudy (anonymous):

that's what I initially thought....

OpenStudy (anonymous):

so now we have to integrate \[=\sum_{0}^{\infty} (-1)^nt^n\]

OpenStudy (anonymous):

yes !

OpenStudy (anonymous):

\[\frac{(-1)^{n+1}}{n+1}\frac{t^{n+1}}{n+1}\]

OpenStudy (anonymous):

no ! \[\sum_{0}^{}(-1)^nt^ {n+1}/n+1 \]

OpenStudy (anonymous):

It was for Mathsophia

OpenStudy (anonymous):

so why is is \[(-1)^{n}\] instead of \[(-1)^{n+1}\]

OpenStudy (anonymous):

in your final answer

OpenStudy (anonymous):

\[ \frac 1 {1+t}= \sum_{n=0}^\infty (-1)^{n} t^n\\ \int_0^x\frac 1 {1+t}dt=\sum_{n=0}^\infty (-1)^{n} \frac {x^{n+1} } { n+1 }=\\ x-\frac{x^2}{2}+\frac{x^3}{3}-\frac{x^4}{4}+\frac{x^5}{5}-\frac{x ^6}{6}+\frac{x^7}{7}+O\left(x^8 \right) \]

OpenStudy (anonymous):

Did you get it?

OpenStudy (anonymous):

My last post is a complete solution. I have to go to bed.

OpenStudy (anonymous):

Thank you Eliassaab!!!

OpenStudy (anonymous):

Yw.

OpenStudy (anonymous):

\[\frac{d}{dx}\sum_{n=0}^\infty(-1)^nx^n\]

OpenStudy (anonymous):

I thought it was the integral and not the derivative? I don't know anymore

OpenStudy (anonymous):

it's the derivative you hve to integrate it

OpenStudy (anonymous):

it is the integral -.-

OpenStudy (anonymous):

gah this is confusing me too since i haven't done this in a while

OpenStudy (anonymous):

its the integral of \[\int \frac{dx}{x+1}\]

OpenStudy (anonymous):

following elissab, the reason why it isn't \[(-1)^{n+1}\frac{x^{n+1}}{n+1}\] is because you are integrating in respects to x and not n

OpenStudy (anonymous):

that is why it is \[(-1)^n\]

OpenStudy (anonymous):

Victory!!!!

OpenStudy (anonymous):

You're awesome!

OpenStudy (anonymous):

I insist that my solution is the correct one. You need to integrate not differentiate. See http://tutorial.math.lamar.edu/Classes/CalcII/PowerSeriesandFunctions.aspx Example 5, where they expand log(5 -x) which is similar to your problem.

OpenStudy (anonymous):

@Outkast3r09 , you are wrong about differentiating. See my post above. Also \[ \frac{d}{dx}\sum_{n=0}^\infty(-1)^nx^n= -\frac{1}{(x+1)^2} \] @experimentX @lalaly @dpaInc @mukushla

OpenStudy (anonymous):

That is the final solution \[ \frac 1 {1+t}= \sum_{n=0}^\infty (-1)^{n} t^n\\ \int_0^x\frac 1 {1+t}dt=\sum_{n=0}^\infty \int_0^x(-1)^{n} t^ndt=\\ \sum_{n=0}^\infty (-1)^{n}\left[\frac{ t^n } {n+1}\right]_0^x= \\ \sum_{n=0}^\infty (-1)^{n} \frac {x^{n+1} } { n+1 }=\\ x-\frac{x^2}{2}+\frac{x^3}{3}-\frac{x^4}{4}+\frac{x^5}{5}-\frac{x ^6}{6}+\frac{x^7}{7}+O\left(x^8 \right) \]

OpenStudy (anonymous):

i go with @eliassaab

OpenStudy (anonymous):

agree with @mukushla ...:)